OpenStudy (anonymous):

How do you find the inverse of logarithmic functions? Ex:h(x) = log7 x Ex:f(x) = 3^x - 1 Ex:h(x) = e^2x Ex:f (x) = 6x

OpenStudy (anonymous):

rewrite in equivalent exponential form \[\log_7(x)=y\iff 7^y=x\]

OpenStudy (s3a):

What satellite73 said, which basically is that you raise the base by the logarithm of the same base.

OpenStudy (anonymous):

So is it the same for all of those?

OpenStudy (s3a):

Ex:h(x) = log7 x THIS IS THE ONLY LOGARITHMIC FUNCTION YOU LISTED. Ex:f(x) = 3^x - 1 THIS IS AN EXPONENTIAL FUNCTION. Ex:h(x) = e^2x THIS IS AN EXPONENTIAL FUNCTION. Ex:f (x) = 6x THIS IS A LINEAR FUNCTION. So, it is the same for all LOGARITHMIC functions.

OpenStudy (anonymous):

The other ones said that the answer was logarithmic

OpenStudy (s3a):

If a function is exponential, you need to take the logarithm of both sides to get the answer of x (as a a logarithmic function).

OpenStudy (anonymous):

It says Write the inverse of g(x) = e^x g^-1(x) = _____. a)log e x b)log x e c)log e y so how would you do that?

OpenStudy (s3a):

log_e(x) = ln(x) g^(-1) (x) = ln(e^x) g^(-1) (x) = x Since ln(x) = log_e(x) the logic is "e raised to what equals e^x? The answer is x."

OpenStudy (s3a):

Oops.

OpenStudy (anonymous):

what?

OpenStudy (s3a):

You switch the x and y and then you solve for y and that's the inverse.

OpenStudy (s3a):

You solve for y using what satellite73 said on the first post.

OpenStudy (s3a):

where y = f(x) or h(x) or whatever.

OpenStudy (s3a):

Here: https://www.youtube.com/watch?v=btmXzOSn1tY

OpenStudy (anonymous):

Okay thank you that's a lot of help. This stuff really confuses me

OpenStudy (s3a):

No problem.